The sum of two numbers is 52.The larger number is 2 less than twice the smaller number.What are the numbers?
Answer by Maths68(1474)

Let x is larger and y is smaller number

Given

The sum of two numbers is 52
x+y=52.......(1)

The larger number is 2 less than twice the smaller number.
x=2y-2.......(2)

Put the value of x from (2) to (1)
x+y=52
2y-2+y=52
3y=52+2
3y=54
3y/3=54/3
y=18
Put the value of y in (2)
x=2y-2
x=2(18)-2
x=36-2
x=34



Larger number is 34
Smaller number is 18